### v1

@@ 2 Technical Overview
$\mathbf{B}_{1}=(\mathbf{B}_{i,j}^{(1)})_{i,j=0}^{m-2},\ \mathbf{B}_{2}=(\mathbf{B}_{i,j}^{(2)})_{i,j=0}^{m-2}\quad$ $\leadsto\quad b(x,y,z)=\sum_{i,j=0}^{m-2}\left(\mathbf{B}_{i,j}^{(1)}\cdot y^{i+1}z^{j+1}+\mathbf{B}_{i,j}^{(2)}\cdot x^{i+1}z^{j+1}\right).$ (2.3)
Since [Coh+05] aim for the best-possible $n^{\omega}$, they invoke the STPP construction (2.2) recursively using the Asymptotic Sum Inequality [Sch81]. Our simple observation is that one can instead compute the polynomial multiplication (2.2) directly using the group’s FFT (the group is $\mathbb{Z}_{m}^{3N}$) to compute $d$ products of $q\times q$ matrices, asymptotically faster than the naïve algorithm $O(d\cdot q^{3})$ (albeit far from the optimal bound on $\omega$ achievable via recursion). This is illustrated in Figure 1.
Given this black-box algorithm for computing $d$ products of $q\times q$ matrices, we can use it to multiply $dq\times dq$ square matrices $\mathbf{A},\mathbf{B}$ using the following idea. Partition the matrices into $d\times d$ blocks of size $q\times q$. Then computing $\mathbf{A}\mathbf{B}$ requires exactly $d^{3}$ block multiplications (the same as in the scalar case, just with blocks), hence it suffices to invoke the black box algorithm $d^{2}$ times (which can be done in parallel, unlike [Coh+05]). The runtime savings (if any) of a single invocation of the black-box algorithm are thus multiplied by a factor of $d^{2}$. This greatly improves the asymptotics, and substitutes the role of recursion in the analysis. The exact analysis of this idea for the aforementioned construction leads to the result in Theorem 1.2.
@@
2.2 Approximate Matrix Multiplication via Convolutions
Our main contribution is extending the Cohn Umans framework to the approximate case, i.e., studying approximate analogues of (TPP). and (STPP). sets in abelian groups. Our main motivation here is to mitigate the asymptotics of Algorithm 1 (depicted in Figure 1) and to take another step toward practical MatMul  algorithms, in settings where inexact computation is tolerable (such as Deep Learning [Ail+25], Databases [LLM23], and statistics [Sar06, CW13]). The idea here is to “compress" the CKSU polynomials (2.2) in a way that allows for faster convolution algorithm, without hurting the combinatorial STPP property too much. We provide three low-degree approximation schemes of the (TPP or STPP) polynomials, leading to Theorem 1.3, Theorem 1.4 and Theorem 1.5.
AMM via Sketching Convolutions (Proof of Theorem 1.3)
As a warmup, we show that combining the exact algorithm (1) from Theorem 1.2 with standard linear Sketching ([Sar06, DKM06, CW13]) of the input matrices in a black-box manner, already improves the best known asymptotic speed-accuracy tradeoff (AMM error) achievable by combinatorial algorithms. The proof is straightforward: We first use a random Gaussian matrix $S\in_{R}\mathcal{N}(0,1/\sqrt{n})^{n\times r}$ to “compress" the number of columns (resp. rows) of the input matrices $\mathbf{A}^{\prime}:=\mathbf{A}S$, $\mathbf{B}^{\prime}:=S^{\top}\mathbf{B}$, yielding two $(n\times r)$ matrices. The JL moment property [CW13] then guarantees that
@@ 2.2 Approximate Matrix Multiplication via Convolutions
It therefore remains to multiply $\mathbf{A}^{\prime}\mathbf{B}^{\prime}$, which using naïve MatMul  requires $O(rn^{2})$ time, recovering the standard linear speed-accuracy AMM error. Our obvious approach is to use Theorem 1.2 (or more precisely, a rectangular version of it, similar to the convolution in Figure 1), to multiply $\mathbf{A}^{\prime}\mathbf{B}^{\prime}$ slightly faster, in $\approx r^{0.9}n^{2}$ time. Note that the composed algorithm reduces the dimensions, and therefore the degree, of the STPP polynomials being convolved. Since this yields a (small) polynomial saving in runtime, we can afford to increase the number of rows in the sketching matrix $S$ to $\approx r^{1.1}$, while still matching the original runtime budget $O(rn^{2})$. By (2.4) this yields an error of $\approx\frac{1}{r^{1.1}}\left\lVert\mathbf{A}\right\rVert_{F}^{2}\left\lVert\mathbf{B}\right\rVert_{F}^{2}$.
@@
We remark that Pagh’s TensorSketch algorithm [Pag13] also uses FFT (albeit in a different way, to efficiently recover the “heavy-hitters" of an outer product $uv^{\top}\in\mathbb{R}^{n\times n}$), but nevertheless his algorithm still yields only a linear dependence on $r$, see (AMM error).
Sparse Fourier Approximation (Proof of Theorem 1.4)
A suboptimal feature of Theorem 1.3 is that it is agnostic to the structure of the CKSU polynomials (2.2) being convolved, and as an artifact, the sketching algorithm still suffers from the asymptotic caveats of the exact algorithm in Theorem 1.2. The main result of this paper (and also its technical centerpiece) is a non-blackbox low-degree approximation scheme for the CKSU polynomials, based on a surprising Fourier-concentration property of these polynomials. This approximation scheme leads to dramatically better AMM tradeoffs for a large class of product distributions ($\mathbf{A}\perp\mathbf{B}$), both in theory and practice (see section 10). Below we explain the main ideas of this low-degree approximation scheme.
Since convolutions are equivalent to elementwise multiplication in the Fourier domain, Parseval’s identity implies that small Fourier coefficients (frequencies) could be dropped without incurring large (squared) error. Moreover, if most of the energy ($\ell_{2}^{2}$ mass) of the Fourier spectrum of $a,b$ happens to lie on low degree frequencies [^4] In the sense that degree of a character is just the sum of the frequency elements $\xi\in G$ defining it, see section 3., then we obtain a (Fourier basis) low-degree approximation for the polynomials $a,b$, whose convolution is not far (in the $\ell_{2}$ sense) from the exact convolution of the CKSU polynomials $a:=P_{\mathbf{A}},b:=P_{\mathbf{B}}$ in (2.2), which recovers $\mathbf{A}\mathbf{B}$. Truncating both polynomials to the same set of $r$ frequencies (Fourier coefficients), resulting in two new polynomials $a^{\prime},b^{\prime}$, means that we can compute $a^{\prime}*b^{\prime}$ using only $r$ multiplications in the Fourier domain. Thus if we can compute the transform on these frequencies fast, we obtain the desired $\tilde{O}(rn^{2})$ time algorithm in Theorem 1.4. Here, the structure of the subset of retained frequencies comes into play.
For general matrices $\mathbf{A},\mathbf{B}$, expecting the polynomials $\widehat{P_{\mathbf{A}}},\widehat{P_{\mathbf{B}}}$ in (2.2) to be concentrated in the Fourier domain, is too good to be true [^5] For example, a polynomial which has exactly one non-zero monomial will have a perfectly uniform spectrum (uniform in magnitude).. Nevertheless, we prove that for random matrices coming from “nice enough" product distributions $\mathbf{A},\mathbf{B}\sim\mathcal{D}_{A}\times\mathcal{D}_{B}$ (where the marginal entries of $\mathcal{D}_{A},\mathcal{D}_{B}$ are i.i.d), the polynomials $\widehat{P_{\mathbf{A}}},\widehat{P_{\mathbf{B}}}$ are highly concentrated, in the sense that most of the spectrum lies on low-degree coefficients. For the polynomials (2.2), these frequencies turn out to have a very nice structure, namely, they are composed of $3$ “slabs" of the group $\mathbb{Z}_{m}^{3}$ , which can be visualized as a lattice, as in Figure 2 below.
Figure 2: An illustration of the Fourier spectrum of $\mathbb{Z}_{m}^{3}$ viewed as a lattice with $m$ points on each axis. The blue region highlights a single slab of width $r$, as used in Theorem 1.4.
As apparent in (2.2), the polynomial $a:=P_{\mathbf{A}}$ is supported on the monomials $X_{1}=\left\{x^{i}y^{j}\mid 0<i,j\leq m-1\right\}$ and $X_{2}=\left\{y^{i}z^{j}\mid 0<i,j\leq m-1\right\}$. The indicator of the support, $\mathbf{1}_{\mathrm{supp}(a)}$, can be written as $\mathbf{1}_{X_{1}}+\mathbf{1}_{X_{2}}$ (because $X_{1}$ and $X_{2}$ are disjoint). These are polynomials obtained by summing all monomials in $X_{1}$ and $X_{2}$ respectively. It is easily shown that the Fourier transform of $\mathbf{1}_{X_{1}}$ at a frequency $\xi\in\mathbb{Z}_{m}^{3}$ (write $\xi=(\xi_{1},\xi_{2},\xi_{3})$ with $\xi_{i}\in[m]$) is just
$\hat{\mathbf{1}}_{X_{1}}(\xi)=m^{-3/2}\times\begin{cases}(m-1)^{2}&\xi_{1}=\xi_{2}=0,\\ -(m-1)&\xi_{1}=0\text{ xor }\xi_{2}=0,\\ 1&\xi_{1},\xi_{2}\not=0.\end{cases}$
The proof of this fact uses the geometry of complex roots of unity, which are the basic building block of Fourier analysis of finite groups (see section 3). A symmetric result holds for $X_{2}$ and for the indicator of $\mathrm{supp}(b)$. Thus almost all of the Fourier mass of $\mathbf{1}_{\mathrm{supp}(a)}$ resides on facets of the cube, where $\xi_{i}=0$ for some coordinate $i$. The choice of the slabs aims to cover these “heavy” facets. This explains the “diminishing returns” dependence of the error bound in Theorem 1.4 on $r$, as most of the energy is already captured by the thinnest slabs (of width $1$). It is also important to note that this specific structure of slabs lends itself to fast computation of the partial transform and its inverse.
To formalize this intuition, we need to bound the expected ratio between the $\ell^{2}$-mass (energy) outside of the retained frequencies, and the total energy product $\|a\|_{2}^{2}\|b\|_{2}^{2}$. The first step of the proof is a concentration bound of the denominator $\|a\|_{2}^{2}\|b\|_{2}^{2}$, which reduces the problem to bounding the ratio of expectations instead. The concentration bound is just an application of Hoeffding’s inequality, since $\|a\|_{2}^{2}$ and $\|b\|_{2}^{2}$ are just sums of i.i.d. random variables. We pay for this with an added exponentially small error term.
Given this reduction, the proof proceeds as follows. Our goal is to compute the algorithm’s expected non-normalized error. The first step is showing that $\mathbb{E}[\hat{a}(\xi)\cdot\overline{\hat{a}(\eta)}]=C_{a}(\xi-\eta)$ for some function $C_{a}$ which depends on the distribution $\mathcal{D}_{A}$ (similarly, for $b$), i.e., that the above expectation is essentially a function of the difference between the frequencies. The functions $C_{a},C_{b}$ are directly determined by the structure of the polynomials (i.e., their support sets), and can be estimated using a mean and variance part, where the variance part is sensitive to $\delta$ and usually overwhelms the mean part.
The second step, is to express the expected error as a sum of the form $\sum_{\delta\in\mathbb{Z}_{m}^{3}}S\cdot C_{a}(\delta)\cdot C_{b}(\delta)\cdot T(\delta)$, where $S=(m-r)^{3}$ is a constant, and $T(\delta)=\sum_{x\in\mathrm{Output}}\omega^{x\cdot\delta}$ (summing over the output coordinates, from which we read off the result). From here, we continue with a straightforward case analysis and estimate the sum. The obtained upper bound is then used to conclude the theorem.
This bound simplifies for $\left\{0,1\right\}$ random Bernoulli matrices and folded Gaussian matrices (where the values distribute as the absolute value of a Gaussian). This is sufficient to deduce a very strong result for $\left\{\pm 1\right\}$ Rademacher and Gaussian matrices, by splitting the positive and negative parts of the matrices (which distribute like Bernoulli matrices), and using the distributivity of the convolution. In other words, observe that if $a=a_{1}\pm a_{2}$, then $\left\lVert a*b\right\rVert_{2}^{2}\leq 2\left\lVert a_{1}*b\right\rVert_{2}^{2}+2\left\lVert a_{2}*b\right\rVert_{2}^{2}$ (and the same for $b$). The formal details are presented in section 9.
@@ 3 Preliminaries
3.1 Fourier Analysis of Abelian Groups
[…]
The group algebra $\mathbb{C}[G]$, can be endowed with an inner product $\left\langle{f,f^{\prime}}\right\rangle=\sum_{g\in G}f(g)\cdot\overline{f^{\prime}(g)}$. The Artin-Wedderburn theorem states that the Fourier transform $\mathcal{F}_{G}$ of $G$ is an isomorphism of algebras that diagonalizes the algebra $\mathbb{C}[G]$ into a direct sum of matrix algebras $\bigoplus_{i}\mathcal{M}_{d_{i}\times d_{i}}(\mathbb{C})$, where $\left\{d_{i}\right\}$ are called the character degrees of $G$. This provides a bound on the tensor rank of the structure tensor of GG [^6] The structure tensor $\mathcal{T}_{G}$ of a group $G$ is a tensor describing the bilinear convolution operator $\mathbb{C}[G]\times\mathbb{C}[G]\to\mathbb{C}[G]$. Formally, we have variables $\left\{X_{g}\right\}_{g\in G},\left\{Y_{g}\right\}_{g\in G},\left\{Z_{g}\right\}_{g\in G}$ and $\mathcal{T}_{G}(X_{g},Y_{h},Z_{f})=1$ if $gh=f$ and 0 otherwise, exactly matching the definition $\delta_{g}*\delta_{h}=\delta_{gh}$. See Appendix B for more details., $\mathcal{T}_{G}$, which is $\sum_{i}d_{i}^{3}$. In case $G$ is abelian, all degrees are $1$, since all linear representations of $G$ are one-dimensional, and so $\bigoplus_{i}\mathcal{M}_{d_{i}\times d_{i}}(\mathbb{C})\cong\mathbb{C}^{\left|G\right|}$, where the multiplication in the latter algebra is just the element-wise product of vectors. In other words, $\mathcal{F}_{G}$ transforms convolution into element-wise multiplication. This is known as the convolution theorem.
Assuming $G$ is abelian, the dual group $\hat{G}$, which is the set of characters (one dimensional representations, i.e., group morphisms $G\to\mathbb{C}^{\times}$), is non-canonically [^7] It requires us to choose a primitive root of unity of order $\left|G\right|$. Our choice will always be $\exp(\frac{2\pi i}{|G|})$. isomorphic to $G$. Thus $\mathbb{C}[G]\cong\mathbb{C}[\hat{G}$], and so the Fourier transform can be viewed as a change of basis, from the $\delta$-functions $\delta_{g}(h)=\mathbf{1}_{g=h}$, to the basis of characters. To ensure it is unitary, we choose the definition of the Fourier transform to be
$f\mapsto(\hat{f}_{\chi})_{\chi\in\hat{G}}=\frac{1}{\sqrt{\left|G\right|}}\left(\left\langle{f,\chi}\right\rangle\right)_{\chi\in\hat{G}}=\left(\frac{1}{\sqrt{\left|G\right|}}\sum_{g\in G}f(g)\chi(g^{-1})\right)_{\chi\in\hat{G}}$.
Under this normalization, we have Parseval’s theorem $\left\lVert f\right\rVert_{2}=\|\hat{f}\|_{2}$, and the convolution theorem is then stated as $\widehat{f*f^{\prime}}=\sqrt{\left|G\right|}\cdot\hat{f}\odot\hat{f}^{\prime}$. The inverse Fourier transform is given by $\hat{f}\mapsto\frac{1}{\sqrt{G}}\left(\sum_{\chi\in\hat{G}}\hat{f}_{\chi}\cdot\chi(g)\right)_{g\in G}$.
By the structure theorem for abelian groups, $G$ can be written as $G=\mathbb{Z}_{n_{1}}\times\ldots\times\mathbb{Z}_{n_{d}}$ for some integers $n_{i},d$. For $n\in\mathbb{N}$, fix $\omega_{n}$ to be a primitive root of unity of order $n$, then $\hat{\mathbb{Z}}_{n}=\left\{\chi_{m}\mid m\in[n]\right\}$ where $\chi_{m}(x)=\omega_{n}^{a\cdot m}$. Moreover, $\hat{G}$ then $\left\{\otimes_{i=1}^{d}\chi_{m_{i}}\mid m_{i}\in[n_{i}]\right\}$, where $\phi\otimes\psi$ is the function $(x,y)\mapsto\phi(x)\cdot\psi(y)$ given that $\phi,\psi$ are functions into $\mathbb{C}$. We define the degree of a character $\chi\in\hat{G}$, given as $\chi=\otimes_{i=1}^{d}\chi_{m_{i}}$, to be $\sum_{i=1}^{d}m_{i}$. The Fourier transform is then explicitly written as
$\hat{f}(\chi_{(m_{1},\ldots,m_{d})})=\frac{1}{\sqrt{\left|G\right|}}\sum_{g\in G}f(g)\cdot\omega_{n_{1}}^{-g_{1}\cdot m_{1}}\cdots\omega_{n_{d}}^{-g_{d}\cdot m_{d}}.$
The tuple $(m_{1},\ldots,m_{d})\in G$ defining $\chi$ is often called the frequency, and so we can drop the $\chi$ notation.
3.2 Paper Organization
In section 4 we present a few constructions we will need for the rest of the paper. In section 5 we present the fast algorithm and its analysis, thus proving Theorem 1.2. In section 6 we present a formal approximation framework based on abelian group convolutions. In section 7 we present the black box Sketch & Solve algorithm, proving Theorem 1.3. In section 8 we present the small abelian group approximation approach, proving Theorem 1.5. In section 9 we present the Fourier sparsification approach behind Theorem 1.4 and prove the theorem. At last, in section 10 we provide some empiric results from our tests, showing supporting evidence for our approach.
4 Constructions
4.1 Vanilla TPP
@@
4.2 CKSU STPP
This construction was first given in [Coh+05] to answer a long-standing question. Set $H=\mathbb{Z}_{m}^{3}$ and denote each copy of $\mathbb{Z}_{m}$ in $H$ by $H_{1},H_{2},H_{3}$ (meaning $H_{1}=\mathrm{Im}(\phi_{1})$ where $\phi_{1}:\mathbb{Z}_{m}\hookrightarrow H$, $x\mapsto(x,0,0)$, and similarly $H_{2},H_{3}$ are defined). Define
@@ 4.2 CKSU STPP
The authors in [Coh+05] have shown that this sets satisfy the (STPP).. It allows to compute $2$ independent products of $(m-1)\times(m-1)$ matrices. Using B.3, we set $G=H^{N}$ for some $N\geq 1$, and by a single convolution in $G$ we compute $2^{N}$ independent products of $(m-1)^{N}\times(m-1)^{N}$ matrices. The STPP sets can be indexed by a binary string $\boldsymbol{\alpha}\in\left\{0,1\right\}^{N}$:
@@
5 Sub-Cubic Matrix Multiplication via Convolutions (Proof of Theorem 1.2)

### v2

@@ 1.2 Technical Overview
$\mathbf{B}_{1}=(\mathbf{B}_{i,j}^{(1)})_{i,j=0}^{m-2},\ \mathbf{B}_{2}=(\mathbf{B}_{i,j}^{(2)})_{i,j=0}^{m-2}\quad$ $\leadsto\quad b(x,y,z)=\sum_{i,j=0}^{m-2}\left(\mathbf{B}_{i,j}^{(1)}\cdot y^{i+1}z^{j+1}+\mathbf{B}_{i,j}^{(2)}\cdot x^{i+1}z^{j+1}\right).$
Since [Coh+05] aim for the best possible bound on $\omega$ (the matrix multiplication exponent), they invoke this algorithm recursively (done implicitly through the Asymptotic Sum Inequality [Sch81]). Our simple observation is that one can instead compute the polynomial multiplication (1.2) directly using the group’s FFT (the group is $\mathbb{Z}_{m}^{3N}$) to compute $d$ products of $q\times q$ matrices, asymptotically faster than the naïve algorithm $O(d\cdot q^{3})$ (albeit far from the optimal bound on $\omega$ achievable via recursion). This is illustrated in Figure 1.
Given this black-box algorithm for computing $d$ products of $q\times q$ matrices, we can use it to multiply $dq\times dq$ square matrices $\mathbf{A},\mathbf{B}$ by partitioning the matrices into $d\times d$ blocks of size $q\times q$, and invoke the algorithm $d^{2}$ times. This amplifies the runtime savings of the algorithm, and substitutes the role of recursion in the analysis, leading to the exponent in Theorem 1.2.
@@
1.2.2 Sketch and Solve Approximation (Theorem 1.3)
Using the algorithm obtained above, we observe that combining it with standard linear sketching techniques ([Sar06, DKM06, CW13]) of the input matrices in a black-box manner, already improves the best known asymptotic speed-accuracy tradeoff (AMM error) achievable by combinatorial algorithms. The proof is straightforward: We first use a random Gaussian matrix $S\in_{R}\mathcal{N}(0,1/\sqrt{n})^{n\times r}$ to “compress" the number of columns (resp. rows) of the input matrices $\mathbf{A}^{\prime}:=\mathbf{A}S$, $\mathbf{B}^{\prime}:=S^{\top}\mathbf{B}$, yielding two $(n\times r)$ matrices. The JL moment property [CW13] then guarantees that
@@
It therefore remains to multiply $\mathbf{A}^{\prime}\mathbf{B}^{\prime}$, which using naïve MatMul  requires $O(rn^{2})$ time, recovering the standard linear speed-accuracy AMM error. Our obvious approach is to use Theorem 1.2 (or more precisely, a rectangular version of it), to multiply $\mathbf{A}^{\prime}\mathbf{B}^{\prime}$ slightly faster, in $\approx r^{0.9}n^{2}$ time. Note that the composed algorithm reduces the dimensions, and therefore the degree, of the STPP polynomials being convolved. Since this yields a (small) polynomial saving in runtime, we can afford to increase the number of rows in the sketching matrix $S$ to $\approx r^{1.1}$, while still matching the original runtime budget $O(rn^{2})$. By (1.3) this yields an error of $\approx\frac{1}{r^{1.1}}\left\lVert\mathbf{A}\right\rVert_{F}^{2}\left\lVert\mathbf{B}\right\rVert_{F}^{2}$.
@@
We remark that Pagh’s TensorSketch algorithm [Pag13] also uses FFT (albeit in a different way, to efficiently recover the “heavy-hitters" of an outer product $uv^{\top}\in\mathbb{R}^{n\times n}$), but nevertheless his algorithm still yields only a linear dependence on $r$, as in (AMM error).
1.2.3 Approximate Triple Product Polynomials
One natural approach, is to relax the definition of the (TPP)., by allowing non-trivial solutions. We define $\rho_{n}(G)$ to be the minimal number of solutions to the equation
$s^{-1}s^{\prime}t^{-1}t^{\prime}u^{-1}u^{\prime}=1_{G}$
over all indexing sets $S,T,U\subset G$ (which have enough structure to allow embedding $n\times n$ matrices). In other words, $\rho_{n}(G)=\min_{(S,T,U)}1_{S^{-1}}*1_{S}*1_{T^{-1}}*1_{T}*1_{U^{-1}}*1_{U}(e_{G})$. We also extend this definition to the (STPP). case. We note that the case of abelian groups, this quantity is minimized by highly structured additive sets, namely arithmetic progressions. In particular, it can be re-written as
$\min_{(S,T,U)}\left\lVert[(S-U)-(S-U)]\cap(T-T)\right\rVert,$
where $\left\lVert\cdot\right\rVert$ denotes multi-set cardinality. It is easy to show that in the group $\mathbb{Z}_{rn^{2}}$, this quantity is minimized when SUS-U is an arithmetic progression, and $T$ is chosen so that the difference set TTT-T avoids the arithmetic progression of SUS-U.
We show that for a family of randomized algorithms, based on convolving the embedded signals in $G$ after adding some random signs and permuting the indexing sets, the error is related to the quantity $\rho_{n}(G)$ in the following way. We show that the normalized error is lower bounded by $\Omega\left(\frac{\rho_{n}(G)-n^{3}}{n^{4}}\right)$, which we also show is tight. The main takeaway of Theorem 1.4 is that $\rho_{n}(G)$ precisely determines the speed-accuracy tradeoff of the randomized algorithm.
Clearly, the size of $\rho_{n}(G)$ is inherently related to the group structure, and one could hope to find some special sets which successfully achieve a very low number of non-trivial solutions, compared to the group size, thus leading to a super-linear tradeoff. In abelian groups, one can transform this question into a finding sumsets which minimize the multi-intersection with an arithmetic progression. However, a nifty Cauchy-Schwarz argument shows that $\rho_{n}(G)\geq\frac{n^{6}}{\left|G\right|}$, implying that abelian groups cannot break the linear tradeoff. We provide a very simple construction, based on arithmetic progressions in $\mathbb{Z}_{rn^{2}}$, (almost) attaining the lower bound.
For non-abelian groups, the lower bound is a bit more subtle, and requires Fourier analytic machinery. We provide a proof – which is directly based on a proof of Gowers for mixing in non-abelian groups (see [Bre14]) – showing that $\rho_{n}(G)\geq\frac{n^{6}}{\left|G\right|}-n^{3}\cdot\sqrt{\frac{\left|G\right|}{d(G)}}$, where $d(G)$ is the smallest degree of a non-linear complex representation of $G$. This is a much weaker lower bound, as it requires $d(G)\approx r^{3}$, assuming $\left|G\right|\leq rn^{2}$, for it to have any meaningful interpretation. The non-abelian case is less interesting for practical algorithms (due to the lack of FFT and the block-diagonal form), but may be of independent interest. This result leaves the case of small $d(G)$ open.
@@ 2 Preliminaries
2.1 Constructions
Vanilla TPP.
@@ 2.1 Constructions
CKSU STPP.
This construction was first given in [Coh+05]. Set $H=\mathbb{Z}_{m}^{3}$ and denote each copy of $\mathbb{Z}_{m}$ in $H$ by $H_{1},H_{2},H_{3}$ (meaning $H_{1}=\mathrm{Im}(\phi_{1})$ where $\phi_{1}:\mathbb{Z}_{m}\hookrightarrow H$, $x\mapsto(x,0,0)$, and similarly $H_{2},H_{3}$ are defined). Define
@@
The authors in [Coh+05] have shown that this sets satisfy the (STPP).. It allows to compute $2$ independent products of $(m-1)\times(m-1)$ matrices. Using B.1, we set $G=H^{N}$ for some $N\geq 1$, and by a single convolution in $G$ we compute $2^{N}$ independent products of $(m-1)^{N}\times(m-1)^{N}$ matrices. The STPP sets can be indexed by a binary string $\boldsymbol{\alpha}\in\left\{0,1\right\}^{N}$:
@@
3 Sub-Cubic Matrix Multiplication via Convolutions